OpenStudy (anonymous):

What is the correct solution to the equation? 5(x – 5) = 15 A. –50 B. –2 C. 8 D. 80

OpenStudy (anonymous):

x = 8

OpenStudy (anonymous):

Are you trying to work any of these out? The problems you have been posting are really simple

OpenStudy (anonymous):

thanks shay

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

shay thats the wrong answer

OpenStudy (anonymous):

First you would use the distributive property and distribute the 5 into the () and then add 25 to both sides and finally divide by 5

OpenStudy (anonymous):

@BukLoon that is answer I got as well

OpenStudy (anonymous):

Simplifying 5(x + -5) = 15 Reorder the terms: 5(-5 + x) = 15 (-5 * 5 + x * 5) = 15 (-25 + 5x) = 15 Solving -25 + 5x = 15 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '25' to each side of the equation. -25 + 25 + 5x = 15 + 25 Combine like terms: -25 + 25 = 0 0 + 5x = 15 + 25 5x = 15 + 25 Combine like terms: 15 + 25 = 40 5x = 40 Divide each side by '5'. x = 8 Simplifying x = 8

OpenStudy (anonymous):

i screwed up in calculation haha :P sorry

OpenStudy (anonymous):

@BukLoon you have the wrong answer

OpenStudy (anonymous):

Why did you reorder the terms? I would leave it as is and simplify...

OpenStudy (anonymous):

just the way i was taught , everyone has their own way same answer

OpenStudy (anonymous):